OpenStudy (anonymous):

(x-6)^2+(y-7)^2=100 How do you solve this I know (6,7) is the center and 100 is the radius

OpenStudy (godlovesme):

the radius is not 100 , it's \[\sqrt{100}\]

OpenStudy (anonymous):

i meant 10^2

OpenStudy (godlovesme):

yes which makes the radius 10 :)

OpenStudy (anonymous):

I'm not sure how to solve this circle equation thing I'm really lost and a little frustrated ): can you maybe walk me through it? pls?

OpenStudy (godlovesme):

sure i'll try :) what do u need help with?

OpenStudy (anonymous):

(x-6)^2+(y-7)^2=100 solving this type of equation

OpenStudy (godlovesme):

what exactly are u solving ? i mean are u trying to figure out the center or what?

OpenStudy (anonymous):

I'm guessing for x,y I don't know what they're asking they just told me to solve that It's complicated i guess...?

OpenStudy (anonymous):

@Godlovesme

OpenStudy (godlovesme):

i think they're asking u to find the center and radius .. is it multiple choice question?

OpenStudy (anonymous):

I guess I know that for sure No it's like my teacher called me and told me i had to answer a few questions, I was confused when she told me to solve that^^ like how exactly? I told her the radius and center point but i don't understand that much. Either ways, she told me i needed to study on the subject /: (kinda hurt my feelings tbh i think i know enough about this....) Can I ask you about another thing? It's also center circle & radius related lol @Godlovesme

OpenStudy (godlovesme):

sure lol

OpenStudy (anonymous):

how do you find the center and radius of the circle for this x2 + y2 +2x − 4y − 20 = 0

OpenStudy (godlovesme):

ahhh im not good at this :/ i just got help about this a couple days ago but i dont remember a single thing so sorry :( @acxbox22 can u help please :)

OpenStudy (anonymous):

it's ok! I'll figure something out

OpenStudy (godlovesme):

good luck and here's my question it might give u a little hint :) http://openstudy.com/study#/updates/55859cafe4b091b59af14ca1